

********************************************************************
UARS WEEKLY STATUS REPORT - 23 APR 2004
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
67% of each day's data amounting to about 856 minutes of solar
observing time per day.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, Mg
II, occultation, solar pointing offset, and high resolution
continuous monitoring scans for periods of UARS real-time telemetry
or onboard tape recording.
One consequence of the decision of the White Sands Operations Center
to replace TDRS-W with TDRS-8 is that portions of SUSIM's daily scans
will be lost for April 23-24. The notification of the change came
so late in the planning process that the SUSIM schedule of solar
scans could not be modified due to UARS operational constraints.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1996039.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS. The MPG attended the UARS TDRS Change Impact
meeting on Wednesday, April 21, 2004.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 10 April 2004 (Orbit 68834, GMT Day 101) through 16
April 2004 (Orbit 68918, GMT Day 107).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 102/1322 68857 POWER OFF
PEM (Z/H) 103/1136 68871 DAY ONLY TO FULLTIME OPS
PEM (Z/H) 105/2115 68907 FULLTIME TO DAY ONLY OPS
The beta angle ranged between the angles of 52.5 degrees to 67.4
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 68 degrees on 17 April 2004 (DOY 108).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.26 to 1.27 degrees C. Additional
battery performance data for each battery on 09 April and 16 April
2004 are provided in Appendix D.
The clock error ranged from 10.5 to –10.3 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
• OUT-OF-LIMITS = 00
• ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
• FOT = 02
• HALOE = 01
• PEM = 02
• SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
• FOT = 00
Zero (0) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UPDATE:
We were just informed on Tuesday of last week, 13 April 04, that TDRS-
8 will be moved to TDW and TDW will be moved to 172. Unfortunately
TDRS-8 does not have MA service available, instead they have SMA.
Currently UARS is capable of using 3 TDRSs. They are TDE, TDW and
275. We take approximately 36-42 events per day utilizing all three
of these TDRSs. The services that we use are SSA and MA. The SSA
events are used for tape dumps ( 1/day ) and OMNI passes ( 1/month).
All other events are MA.
Option 1 - Get SMA to work
1) SMA service has never been tested on UARS because the service was
never available on a TDRS that we could use.
2) The UPS that we use to schedule the contacts does not have any of
the configurations codes to support SMA events. (In process of being
corrected)
3) Once SMA supports were scheduled on the UPS and transferred to the
CMS, the CMS process hung. (Believe this to be a software issue with
the CMS and will require software fix, regression testing, and
transition to new code)
Option 2 - Use the old TDW at 172
1) FDF would have to give us 172 view periods and rename them TDW
2) FDF would have to give us 172 ephemeris and rename it to TDW or
have FORMATS rename the file for them
3) 172 will have to be added to the UPS database
4) 172 events would have to be scheduled on the UPS and transferred
to CMS. Once on the CMS the FOT would have to manually edit each
forward, return and tracking service from 172 to TDW. (up to 45 hand
edits per day)
Option 3 - Schedule only TDE and 275 events
1) This puts us back to a 2-TDRS mode. This means that all the gain
in realtime data received by going to 3-TDRS will be eliminated.
None of the options are very attractive, but this is where we stand.
During the investigation of the attitude data seen by HRDI the FOT
uncovered an issue with the onboard solar ephemeris. The onboard
solar ephemeris was last updated during the UARS epoch change in late
1996. This data was optimized for a period of six years, which ended
in 2003. The FOT sent a day’s worth of solar ephemeris data to FDF to
compare to their solar ephemeris reference and found a difference of
200-300 arcseconds. A meeting is scheduled with FDF for 20 April 2004
to discuss this error and determine the best method to eliminate it.
Any update to the onboard ephemeris will require simulator testing
before being uplinked to the spacecraft.
MEETINGS
The FOT attended the monthly Eclipse DRB meeting on 12 April 2002.
The purpose of this meeting was to discuss the status of open DRs and
discuss the date of the next build.
The FOT also attended a telecon with the HRDI instrument team to
discuss the attitude error seen in their data and the solar ephemeris
findings.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 07 May 2004 (DOY 128). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments:
Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 16 April 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
68847 101/2127 04-109 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
68857 102/1322 04-110 HALOE POWER OFF
68871 103/1136 04-111 PEM (Z/H) DAY ONLY TO FULLTIME
68882 104/0448 04-112 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
68907 105/2115 04-113 PEM (Z/H) FULLTIME TO DAY ONLY OPS
68918 106/1530 04-114 SUSIM EXCEPTION CLEAR CMD
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 16 April 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 16 April 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
68847 101/2130 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
68882 104/0448 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 16 April 2004
09 April 2004 - GMT Day 100 Beta = 47.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-044.8 +4.82/+3.87 +25.6/-14.4 25.8 1.03 19.6
3 +11.2/-123.2 +3.87/+2.60 +24.8/-14.4 25.8 1.02 19.2
16 April 2004 - GMT Day 107 Beta = 67.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-112.0 +3.87/+2.60 +18.8/-14.4 26.4 1.05 15.5
3 +00.0/-050.4 +2.91/+1.65 +18.4/-14.4 26.4 1.05 15.1
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 16 April 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-017 FOT 68744 095/0144 35517 Data drop out/11 minutes
04-018 FOT 68752 095/1342 N/A COP to INIT during commanding
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.